How do you solve:
8.1 + 3.8h - 5.6h = -7.2

Answer by stanbon(75887)

solve:
8.1 + 3.8h - 5.6h = -7.2
Multiply thru by 10 to get rid of the decimals:
81 + 38h - 56h = -72
-18h = -153
h = 8.5
===============
Cheers,
Stan H.
===========